Previous research shows that 20 percent of adults who drink non-diet cola prefer Coca-Cola to Pepsi. Recently, an independent research firm questioned a random sample of 10 adult non-diet cola drinkers.     

a) The chance that exactly 6 of these people will prefer Coca-Cola is:

b) The chance that 6 or more of these people will prefer Coca-Cola is:

c) Calculate mean and standard deviation
Answer #1

Sololion a) Here have to find p(x=6) we Given n-10 and- p = 0.20 q=l-p=0.80. using Binomial distribution plx =o) = 10 (0.20)